proof of cotAcotB+cotBcotC+cotCcotA=1

Grade:

1 Answers

Sher Mohammad IIT Delhi
9 years ago
A + B = 180-C

take tan of both sides

(tan A + tan B)/(1-tan A tan B) = - tan C

or tan A + tan B + tan C = tan A tan B tan C

divide both sides Tan A tan B tan C

1/tan B 1/ tan C + 1/tan A 1/ tan C + 1/ tan A 1/tan B = 1

cot B cot C+ cot A cot C + cot A cot B = 1

sher mohammad
b.tech, iit delhi